What is the cost of capital for the following?

Bond Issue

$6,000,000

30-year maturity

10% coupon rate at 98% of face value

2% of face value would be charged as an underwriting commission
Answer #1

HERE PAYMENT FREQUENCY IS NOT GIVEN
CASE 1: ASSUMING ANNUAL COUPON PAYMENTS
Using financial calculator
N=30*1
PMT=10%*100/1=10
FV=100
PV=-98+2%*100=-96
CPT I/Y=10.4400%

Cost of capital is 10.440%

CASE 2: ASSUMING SEMI-ANNUAL COUPON PAYMENTS
Using financial calculator
N=30*2=60
PMT=10%*100/2=5
FV=100
PV=-98+2%*100=-96
CPT I/Y=5.2191%

Cost of capital is 5.2191%*2=10.4382%